OpenStudy (anonymous):

Evaluate the integral with Trig substitution: integral of x/sqrt((x^2)+x+1) dx

OpenStudy (irishboy123):

\(\int \dfrac{x}{\sqrt{x^2+x+1}} dx\) ?? thought about completing the square in the denominator? \(\int \dfrac{x}{\sqrt{(x+\frac{1}{2})^2+\frac{3}{4}}} dx\) ?? then something like u = x + 1/2 ??

OpenStudy (anonymous):

Yeah I've did that, the part I get stuck on is the part after that

OpenStudy (anonymous):

I've simplified it to say \[\int\limits (2u-1)/\sqrt (4u^2+3)\]

OpenStudy (irishboy123):

well, if \(\int \dfrac{x}{\sqrt{(x+\frac{1}{2})^2+\frac{3}{4}}} dx\) works, i'd go \( u = x + \frac{1}{2}, \quad du = dx\) that leads to \(\int \dfrac{u - \frac{1}{2}}{\sqrt{u^2+\frac{3}{4}}} du\) \(= \int \dfrac{u}{\sqrt{u^2+\frac{3}{4}}} du - \int \dfrac{\frac{1}{2}}{\sqrt{u^2+\frac{3}{4}}} du\) **if** the algebra is correct, that is now only a few lines away from completion :p **if** !!

OpenStudy (anonymous):

okay yeah, I got the correct answer, thank you!